OpenStudy (anonymous):

Find cos θ if csc θ = 2 and θ is in quadrant II.

OpenStudy (anonymous):

OK so... Sin θ = 1/Csc θ so it would be 1/2 which is O/H? So the mesurements can be calculated using x^2 + y^2 = r^2?

OpenStudy (anonymous):

θ=60

OpenStudy (anonymous):

as......... if ....secθ=2 hence ...1/cosθ=2 cosθ=1/2 sin(90+θ)=1/2

OpenStudy (anonymous):

hence .......θ=60

OpenStudy (dumbcow):

csc = 1/sin --> 1/sin = 2 --> sin = 1/2 cos^2 = 1-sin^2 cos^2 = 1-(1/2)^2 cos^2 = 3/4 cos = sqrt3/2

OpenStudy (anonymous):

thought since sin would = 1/2 and csc = 2 then wouldn't it be 30?

OpenStudy (dumbcow):

no theta = 150 since in quad 2

OpenStudy (anonymous):

of it's cosec θ then dumbcow is right

OpenStudy (anonymous):

yea sqrt(3)/2

OpenStudy (anonymous):

would it be negative since it's in the II quadrant? and only thing positive in II quadrant is Sin and CSC?

OpenStudy (anonymous):

sry fakshon ...i consideed it as sec θ

OpenStudy (dumbcow):

correct, cos would be negative :)

OpenStudy (dumbcow):

@shivamsinha it happens :)

OpenStudy (anonymous):

yep yep it's all good.

OpenStudy (anonymous):

thanks for the help.